OpenStudy (anonymous):

To paint his house, Samuel bought 2 cans of paint and one roller for 30$. When friends came over to help, he went back to the store and bought 3 more cans of paint and 2 more rollers for 50$. How much did he pay for each roller?

OpenStudy (anonymous):

2 cans + 1 roller = $30 3 cans + 2 rollers = $50 call the cans C and the rollers R solve for R

OpenStudy (anonymous):

2C+1R=$30 3C+2R=$50

OpenStudy (anonymous):

C+R=$20